OpenStudy (alante):

What is the solution of the system? -4x + 3y = -16 -x + 2y = -4 A. (4, 0) B. (0, 4) C. No solution D. All real numbers

MsBrains (ms-brains):

Use Cymath

OpenStudy (jjcoolja12):

Problem: Solve −4x+3y=−16;−x+2y=−4 Steps: I will try to solve your system of equations. −x+2y=−4;−4x+3y=−16 Step: Solve−x+2y=−4for x: −x+2y+−2y=−4+−2y(Add -2y to both sides) −x=−2y−4 −x −1 = −2y−4 −1 (Divide both sides by -1) x=2y+4 Step: Substitute2y+4forxin−4x+3y=−16: −4x+3y=−16 −4(2y+4)+3y=−16 −5y−16=−16(Simplify both sides of the equation) −5y−16+16=−16+16(Add 16 to both sides) −5y=0 −5y −5 = 0 −5 (Divide both sides by -5) y=0 Step: Substitute0foryinx=2y+4: x=2y+4 x=(2)(0)+4 x=4(Simplify both sides of the equation) Answer: x=4 and y=0

OpenStudy (jjcoolja12):

so its a